Two blocks are sliding across a horizontal frictionless tabletop. Block A has a mass of 5 kg, and it is sliding directly southward with a speed of 8 m/s. Block B has a mass of 7 kg, and it is sliding directly westward with a speed of 4 m/s. The blocks then collide and stick together. What is the speed of the new combined block after the collision?
Answer #1

conservation of momentum

m1 v1 + m2 v2 = (m1 + m2) v

5* 8 (-j) - 7* 4 ( i) = ( 5+7) v

v = - 2.33 i - 3.33 j

net velocity

v^2 = 2.33^2 + 3.33^2

v = 4.067 m/s

====

Comment before rate in case any doubt, will reply for sure.. goodluck